The supplement of an angle is 26 more than 5 times its complement. Find the measure of the angle.
Found 2 solutions by richwmiller, JimboP1977:
Answer by richwmiller(17219)

s 180
c 90
a+s=180
a+c=90
s=26+5c
a=74 c=16 s=106

Answer by JimboP1977(311)

Supplement = 180-angle
Complement = 90-angle
(180-angle)-26=5(90-angle)
154-angle=450-5*angle
-angle=296-5*angle
4*angle=296
angle = 74 degrees.